OpenStudy (anonymous):

Integral of cost/(1+(sint)^2

OpenStudy (anonymous):

\[let~i=\int\limits \frac{ \cos t }{ \left( 1+\sin t \right)^{2} }dt\] put 1+sin t=u cos t dt=du \[I=\int\limits u ^{-2}du=\frac{ u ^{-1} }{-1 }+c=\frac{ -1 }{ u }+c\] put the value of u and get the solution.